# Gas meets Kozai: the influence of a gas-rich accretion disc on hierarchical triples undergoing von Zeipel-Lidov-Kozai oscillations

### v1

@@ 4.1 Inner Orbital Evolution: Transition between Inspiral and Outspiral
[…]
@@
Figure 5: Same as Fig. 4 but comparing a stellar binary with its radii increased by a factor of 3×3$\times$3 × (blue) to a black hole binary (green), for which $A_{\rm geom}=0$. The dependence of the direction of orbital evolution (hardening or softening) on the regime of drag is clearly evident by comparing the fourth row of panels.
From the remaining $\propto v^{2}$ terms, we see from Eq. (25) that inspiral should always occur in the geometric drag regime. This is corroborated by the blue lines shown in Fig. 5, which shows the evolution of a stellar binary with artificially inflated (3×3$\times$3 ×) radii [^2] Physically, this is analogous to a decrease in $a_{\rm out}$ (Eq. 12), but we have chosen to directly modify $R$ so that the ZLK evolution is unaffected, facilitating easier comparison among evolutionary tracks.. On the other hand, for the BHL drag regime, we see that both in and outspiral can occur. This can be seen from the green lines Fig. 5, where we show that a binary black hole undergoing ZLK experiences consistent orbital expansion.
@@
In summary, the combination of the drag transitioning to the BHL regime, as well as the inclination variations of the inner orbit (due to the ZLK effect), causes binaries to widen as they align with the disc. Such a widening weakens the short-range forces experienced by the orbit (Eqs. 4, 6), allowing the system to reach smaller pericentre distances. This has important astrophysical applications as it can potentially induce stellar mergers, as can be seen in bottom-right panel of Fig. 3 (cf. Stephan et al., 2016). We further discuss the consequences of this evolution in Section 5.
@@ 5.3 Caveats & Future directions
First, the interaction of objects with gas-rich environments is complex and not well understood. For simplicity, we have adopted here the simple prescription accounting for the physical and accreting radii of the binary components (Eq. 10). Most notably, we have neglected the effect of gas dynamical friction (Ostriker, 1999). However, because the physical scalings for dynamical friction are the same as that of BHL accretion-driven drag (e.g., Wang et al., 2024), the inclusion of gas dynamical friction is not expected to change our key results concerning the boundary between the hardening and softening of the inner binary, only the rate at which it occurs. The gas can also directly affect the binary components when there is significant mass accretion (Spieksma & Cannizzaro, 2025). In addition, the binaries considered here are soft: their orbital velocities are smaller than the typical velocity dispersion of other stars in the nuclear star cluster. As such, two-body relaxation will also act to gradually soften the binary (Heggie, 1975; Collins & Sari, 2008; Winter-Granic et al., 2024).
@@
Our finding that binaries tend to transition from orbital hardening to softening as their outer orbits align with the disc must be qualified: as the inclination approaches $H/r$, some of the approximations made break down: the gas can follow the binary as $v_{\rm rel}\to c_{s}$, invalidating the $\sim$ ballistic approximation we’ve implicitly assumed by treating the disc as a homogeneous medium with density $\rho_{g}$. Additionally, $\boldsymbol{\mathbf{v}}_{\rm out}$ will no longer be approximately constant as the arclength of the orbit within the disc becomes non-negligible.
@@ Acknowledgements
We thank Mark Dodici, Jeremy Goodman, Melaine Saillenfest, and Scott Tremaine for helpful comments that improved the quality of this manuscript. YS is supported by a Lyman Spitzer, Jr. Postdoctoral Fellowship at Princeton University. The research leading to this work was supported by the Independent Research Fund Denmark via grant ID 10.46540/3103-00205B.
@@ A.1 Secular Equations
×$\times\left[\frac{k_{2,1}m_{1}}{m_{2}}\left(\frac{R_{1}}{a}\right)^{5}+\frac{k_{2,2}m_{2}}{m_{1}}\left(\frac{R_{2}}{a}\right)^{5}\right]n,$ (39)
@@ A.3 Precession due to Disc
We aim to evaluate the nodal precession rate of a particle with semi-major axis $a$ due to a razor-thin disc with inner and outer radii $r_{\rm in}<a<r_{\rm out}$. As discussed in the main text, the standard secular approaches are inapplicable due to the divergence when $a=r$. Instead, we estimate the precession frequency by computing the various epicyclic frequencies introduced by the disc’s gravitational potential: recall that the closed nature of the Keplerian orbit is due to the equality of the radial and vertical epicyclic frequencies with the orbital frequency, and deviations give rise to precession. We specialize to a $p=-1$ surface density profile ($\Sigma=\Sigma_{0}a_{0}/a$), which is known as a Mestel disc (Binney & Tremaine, 2008), and consider circular orbits for simplicity. The potential of such a thin disc can be analytically expressed cylindrical coordinates as (Binney & Tremaine, 2008)
@@
where $J_{0}$ is a Bessel function of the first kind. The vertical acceleration at height $z$ is given by
$\ddot{z}$ $=-\frac{\partial\Phi}{\partial z}$
$=-2\pi G\Sigma_{0}a_{0}\int\limits_{0}^{\infty}J_{0}(ka)e^{-kz}\;\mathrm{d}k$
$=-2\pi G\frac{\Sigma_{0}a_{0}}{a}\frac{1}{(1+(z/a)^{2})^{1/2}},$ (53)
@@
From Eq. (53), we see that a particle at rest at height $z_{0}$ (corresponding to an orbit with inclination $\tan i=z_{0}/a$) will experience (anharmonic) vertical oscillations with period $4\sqrt{2z_{0}/|\ddot{z}|}$. This results in an effective vertical epicyclic frequency (which drives nodal regression, or retrograde orbital precession) of
$\kappa_{z}$ $=\frac{\pi}{2}\sqrt{\pi G\Sigma\frac{1}{a}\cot i\cos i}$
$\Omega_{\rm out}$ $\sim\sqrt{\pi G\rho_{g}(H/a)\cot i\cos i}\frac{\pi}{2}.$ (55)
Note that $\kappa_{z}$ formally diverges as $i\to 0$. This is because $\ddot{z}\neq 0$ when $z=0$. Of course, in three-dimensional discs, the disc cannot be treated as razor-thin when $i\lesssim H/r$ and the orbit is embedded. Since we do not consider low-inclination orbits in this paper, Eq. (55) is appropriate. In practice, we adopt a constant, effective $g_{\rm eff}$ satisfying $g\ll g_{\rm eff}\ll 2\pi/t_{\rm ZLK}$, which gives the qualitatively correct evolution without needing to resolve the rapid orbital precession.
@@ Appendix B Eccentricity Evolution of Outer Binary
The eccentricity evolution of the outer binary is in general a function of the surface density profile of the disc, the drag prescription, and the argument of periastron of the outer orbit ($\omega_{\rm out}$). Results have been derived for specific surface density profiles (e.g., Wang et al., 2024) and for a static gas background (Szölgyén et al., 2022). Here, we provide a simple derivation of the boundary between eccentricity pumping and damping. For simplicity, we assume just a single object (and drop the ‘out’ subscripts). For now, we take $\omega=0$; eccentricities tend to favor damping as $\omega$ is taken to be further from 00 and $\pi$ (Wang et al., 2024), so our result should be taken as a necessary but not sufficient requirement for eccentricity pumping.

### v2

@@ 4.1 Inner Orbital Evolution: Transition between Inspiral and Outspiral
$\approx{}$ $\mu\Bigg{(}\frac{\boldsymbol{\mathbf{F}}_{\rm d,2}\left(\boldsymbol{\mathbf{v}}_{\rm out,rel}\right)+\delta\boldsymbol{\mathbf{F}}_{\rm d,2}\left(\boldsymbol{\mathbf{v}}_{\rm 2}\right)}{m_{2}}$
@@
Figure 5: Same as Fig. 4 but comparing a stellar binary with its radii increased by a factor of $3\times$ to a black hole binary ($R=0$), for which $A_{\rm geom}=0$. The dependence of the direction of orbital evolution (hardening or softening) on the regime of drag is clearly evident by comparing the fourth row of panels.
From the remaining $\propto v^{2}$ terms, we see from Eq. (25) that inspiral should always occur in the geometric drag regime. This is corroborated by the blue lines shown in Fig. 5, which shows the evolution of a stellar binary with artificially inflated ($3\times$) radii [^4] Physically, this is analogous to a decrease in $a_{\rm out}$ (Eq. 12), but we have chosen to directly modify $R$ so that the ZLK evolution is unaffected, facilitating easier comparison among evolutionary tracks.. On the other hand, for the BHL drag regime, we see that both in and outspiral can occur. This can be seen from the green lines Fig. 5, where we show that a binary black hole undergoing ZLK experiences consistent orbital expansion.
@@
In summary, the combination of the drag transitioning to the BHL regime, as well as the orientation of the inner orbit when experiencing the ZLK effect, causes binaries to widen as they align with the disc. Such a widening weakens the short-range forces experienced by the orbit (Eqs. 4, 6), allowing the system to reach smaller pericentre distances. This has important astrophysical applications as it can potentially induce stellar mergers, as can be seen in bottom-right panel of Fig. 3 (cf. Stephan et al., 2016). We further discuss the consequences of this evolution in Section 5.
@@ 5.3 Caveats & Future directions
First, the interaction of objects with gas-rich environments is complex and not well understood. For simplicity, we have adopted here the simple prescription accounting for the physical and accreting radii of the binary components (Eq. 10). Most notably, we have neglected the effect of gas dynamical friction (Ostriker, 1999). However, because the physical scalings for dynamical friction are the same as that of BHL accretion-driven drag (e.g., Wang et al., 2024), the inclusion of gas dynamical friction is not expected to change our key results concerning the boundary between the hardening and softening of the inner binary, only the rate at which it occurs. Notably, such a transition also appears when explicitly modeling the competition between gas dynamical friction and accretion (Grishin & Perets, 2016), and the analytical results presented in this paper can in principle be extended to predict the transition between in- and out-spiral in the presence of gas dynamical friction as well. The gas can also directly affect the binary components when there is significant mass accretion (Spieksma & Cannizzaro, 2025). In addition, the binaries considered here are soft: their orbital velocities are smaller than the typical velocity dispersion of other stars in the nuclear star cluster. As such, two-body relaxation will also act to gradually soften the binary (Heggie, 1975; Collins & Sari, 2008; Winter-Granic et al., 2024).
@@
Our finding that binaries tend to transition from orbital hardening to softening as their outer orbits align with the disc must be qualified: as the inclination approaches $h\equiv H/r$, some of the approximations made break down: the gas can follow the binary as $v_{\rm rel}\to c_{s}$, invalidating the $\sim$ ballistic approximation we’ve implicitly assumed by treating the disc as a homogeneous medium with density $\rho_{g}$. Additionally, $\boldsymbol{\mathbf{v}}_{\rm out}$ will no longer be approximately constant as the arclength of the orbit within the disc becomes non-negligible.
@@
In accretion discs, gas exhibits differential shear motion, which deviates from Keplerian motion. In this paper, we ignore this effect and assume Keplerian velocities. Including corrections due to shearing motion could add corrections to the results derived above, and we leave this for future work.
@@ Acknowledgements
We thank the anonymous referee for helpful comments that improved the quality of this paper. We thank Mark Dodici, Jeremy Goodman, Melaine Saillenfest, and Scott Tremaine for helpful comments and suggestions that improved the quality of this manuscript. This work was inspired by conversations at the New Ideas on the Origin of Black Hole Mergers workshop at the Niels Bohr Institute. YS is supported by a Lyman Spitzer, Jr. Postdoctoral Fellowship at Princeton University. The research leading to this work was supported by the Independent Research Fund Denmark via grant ID 10.46540/3103-00205B.
@@ A.1 Secular Equations
$\times\left[\frac{k_{2,1}m_{1}}{m_{2}}\left(\frac{R_{1}}{a}\right)^{5}+\frac{k_{2,2}m_{2}}{m_{1}}\left(\frac{R_{2}}{a}\right)^{5}\right]n,$ (39)
@@ A.3 Precession due to Disc
In this appendix, we provide a simple derivation for the nodal precession rate of a particle with semi-major axis $a$ due to a razor-thin disc with inner and outer radii $r_{\rm in}<a<r_{\rm out}$. As discussed in the main text, standard secular approaches are inapplicable due to the divergence when $a=r$. Instead, we estimate the precession frequency by computing the various epicyclic frequencies introduced by the disc’s gravitational potential: recall that the closed nature of the Keplerian orbit is due to the equality of the radial and vertical epicyclic frequencies with the orbital frequency, and deviations give rise to precession. For simplicity, we specialize to a p=1p=-1 surface density profile ($\Sigma=\Sigma_{0}a_{0}/a$), which is known as a Mestel disc (Binney & Tremaine, 2008), and focus on circular orbits. The potential of such a thin disc can be analytically expressed in cylindrical coordinates as (Binney & Tremaine, 2008)
@@
where $J_{0}$ is a Bessel function of the first kind. The resulting vertical acceleration at height $z$ due to the disk is given by
$\ddot{z}_{\rm disk}$ $=-\frac{\partial\Phi}{\partial z}$
$=-2\pi G\Sigma_{0}a_{0}\operatorname{sgn}(z)\int\limits_{0}^{\infty}J_{0}(ka)e^{-kz}\;\mathrm{d}k$
$=-2\pi G\frac{\Sigma_{0}a_{0}}{a}\operatorname{sgn}(z)\frac{1}{(1+(z/a)^{2})^{1/2}}$
$=-2\pi G\frac{\Sigma_{0}a_{0}}{a}\operatorname{sgn}(z),$ (53)
@@
This additional anharmonic vertical acceleration adds to the Keplerian vertical restoring force. The resulting vertical epicyclic frequency is given by seeking solutions of form $z=A\sin(\omega t)$ to the differential equation
$\ddot{z}+n^{2}z=-2\pi G\frac{\Sigma_{0}a_{0}}{a}\operatorname{sgn}(z).$ (55)
The dominant oscillation frequency ($\omega\approx n$, as long as $G\Sigma_{0}a_{0}/a\ll n$) can be used to compute the precession frequency $\Omega_{\rm out}\equiv n-\omega$, resulting in
$\Omega_{\rm out}$ $\simeq-\frac{4G}{nz_{0}}\frac{\Sigma_{0}a_{0}}{a}$
$=-\frac{G\rho_{g}}{n}\frac{4h}{\sin i},$ (56)
where $h\equiv H/r$, and the negative sign indicates retrograde precession. Note that $\Omega_{\rm out}$ diverges as $i\to 0$, since $\ddot{z}_{\rm disk}\neq 0$ as $z\to 0$. Physically, this is regularized when $i\lesssim h$, where the orbit is embedded in the the disc. Since we do not consider low-inclination orbits in this paper, Eq. (56) is appropriate. Nevertheless, we note that this closely resembles the standard expression of $\Omega_{\rm out}\simeq\pi G\rho_{g}/n$ for orbits embedded in the disk (i.e., $\sin i<h$; Ward, 1981; Hahn, 2003; Zanazzi & Lai, 2018), suggesting a natural transition between the two regimes.
@@ Appendix B Eccentricity Evolution of Outer Binary
The eccentricity evolution of the outer binary is in general a function of the surface density profile of the disc, the drag prescription, and the argument of periastron of the outer orbit ($\omega_{\rm out}$). Results have been derived for specific surface density profiles (e.g., Wang et al., 2024) and for a static gas background (Szölgyén et al., 2022). Here, we provide a simple derivation of the boundary between eccentricity pumping and damping. For simplicity, we assume just a single object (and drop the ‘out’ subscripts). For now, we take $\omega=0$; eccentricities tend to favor damping as $\omega$ is taken to be further from 0 and $\pi$ (Wang et al., 2024), so our result should be taken as a necessary but not sufficient requirement for eccentricity pumping.